OpenStudy (anonymous):

find all the rational zeros of the function f(x)=6x^4-18x^3-318x^2-54x

OpenStudy (anonymous):

0 is obivious

OpenStudy (anonymous):

since we took zero out , lets divide by (x-0)=x 6x^3 -18x^2-318x-54

OpenStudy (anonymous):

Okay so then?

OpenStudy (anonymous):

Im sorry I just do not get math at all

OpenStudy (anonymous):

divide every factor of 54 by every factor of 6 \[\frac{\pm1,2,3,6,9,27}{1,2,3,6}\] so try 9 6(9)^3 -18(9)^2-318(9)-54=0 yes

OpenStudy (anonymous):

so it would be (0,9)?

OpenStudy (anonymous):

answers are 0 and9

OpenStudy (anonymous):

Okay thank you :)